I have another problem to check for my daughter
factoring SPECIAL CASE expressions
9w^2+60w+100
I don't even no where to begin, please help me again. THANK YOU!!!

Answer by josmiceli(19441)

9w^2+60w+100
first I look at 9w^2 and I see that 9w^2 = (3w)*(3w)
so I start to set up the factors
(3w + ?)(3w + ?)
If I pick 10 * 10 as the factors of 100, the "inner" product is 60w
(3w + 10)(3w + 10) = 9w^2+60w+100
of course,
(3w + 10)(3w + 10) = (3w + 10)^2